# Linking Potentially Misclassified Healthy Food Access to Diabetes Prevalence

### v1

@@ 4.1 Data and Objectives
We investigated the associations between having access to healthy foods and diabetes prevalence across $N=387$ census tracts in the Piedmont Triad region of northwestern North Carolina. A census tract is a small, relatively permanent statistical subdivision of a county or statistically equivalent entity (United States Census Bureau, 2022). Approximately 1.7 million people live in the Piedmont Triad, and the region is known to have excellent transportation infrastructure. The model of interest is defined as $\log\left\{\textrm{E}_{\boldsymbol{\beta}}(Y\mid X,Z)\right\}=\beta_{0}+\beta_{1}X+\beta_{2}Z+\beta_{3}X\times Z+\log(O)$, where $Y$ is the number of diabetes cases, $O$ is the population (used as an offset), $X$ is an indicator of having access to healthy foods within a given radius, and $Z$ is an indicator of metropolitan status. We calculated the Haversine and route-based distances from the population-weighted centroid, according to the 2020 U.S. Census, to the nearest healthy food retailer. Using one-mile and half-mile radii, we discretized these distances to create our exposures $X$ and $X^{*}$. For example, $X=1$ if a census tract has at least one healthy food retailer within the threshold and $=0$ otherwise.
Queried data (i.e., with both $X$ and $X^{*}$) were available for all census tracts. However, to demonstrate the two-phase design, we assumed that only a subset had available route-based access data for the MLE and complete case analyses. After running the gold standard analysis on the full data, we created a partially queried version wherein $X$ was only available for census tracts with $X^{*}=0$ (such that $X=0$ necessarily) and a subset of those with $X^{*}=1$. A random sample of $n=77$ tracts with $X^{*}=1$, targeting a 50/50 split between metropolitan/non-metropolitan, was chosen. Only tracts with error-prone access (i.e., $X^{*}=1$) needed to be validated, since those with $X^{*}=0$ were already known to have $X=0$ and do not inform the misclassification mechanism. The naive analysis used the error-prone access variable on the entire sample, and the complete case analysis used error-free access variable on the subset of queried data. The MLE used the error-prone and error-free access indicators when available and just the error-prone version otherwise.
@@ 4.2 Diabetes Prevalence and Healthy Food Access
At both half- and one-mile radii, the majority of tracts in the Piedmont Triad ($n=353$ [$91\%$] and $n=265$ [$68\%$], respectively) did not have access to healthy foods (Figure 1). Larger, rural tracts tended not to have access. Smaller, urban tracts tended to have access, or at least to be misclassified as such based on Haversine distances. Of the $192$ tracts with access to healthy foods based on one-mile Haversine distances, only $122$ actually had it based on route-based ones ($PPV=122/192=64\%$). As expected, there was more misclassification at the smaller threshold ($PPV=34/85=40\%$).
@@
The estimated prevalence ratio $\widehat{PR}_{non}=\exp(\hat{\beta_{1}})$ compared the predicted prevalence of diabetes in non-metropolitan census tracts with versus without access to healthy foods. In metropolitan census tracts, $\widehat{PR}_{metro}=\exp(\hat{\beta_{1}}+\hat{\beta_{3}})$ was interpreted similarly, and 95% CIs for this prevalence ratio were calculated as $\exp\left\{(\hat{\beta_{1}}+\hat{\beta_{3}})\pm 1.96\widehat{\textrm{SE}}(\hat{\beta_{1}}+\hat{\beta_{3}})\right\}$, where $\widehat{\textrm{SE}}(\hat{\beta_{1}}+\hat{\beta_{3}})=\sqrt{\widehat{\textrm{V}}(\hat{\beta_{1}})+\widehat{\textrm{V}}(\hat{\beta_{3}})+\widehat{\textrm{Cov}}(\hat{\beta_{1}},\hat{\beta_{3}})}$. We focus on these two prevalence ratios; the full models are summarized in Supplemental Table S4.
At either threshold, prevalence of diabetes in neighborhoods with access to healthy foods was expected to be higher than in those without (Figure 2). This counterintuitive trend held at both distance thresholds and across all analysis methods. For example, based on the gold standard, diabetes prevalence in metropolitan tracts with access to healthy foods at one mile was expected to be 18% higher than in those without ($\widehat{PR}_{metro}=1.18$; 95% CI: $1.16$, $1.21$); among non-metropolitan tracts, those with access were expected to have 9% higher prevalence ($\widehat{PR}_{non}=1.09$; 95% CI: $1.06$, $1.11$).
@@
The difference between tracts with and without access was more pronounced in metropolitan areas, according to the gold standard analyses. At both thresholds, the MLE also captured this difference, while the complete case analysis only did so at the one-half mile radius, and the naive analysis did not catch it at either threshold. The MLE actually exaggerated this difference at the one-mile threshold but otherwise closely aligned with the gold standard. Its 95% confidence intervals were slightly narrower than the complete case. For more details, see Web Appendix C.
@@ 5 Conclusion
We introduced a novel maximum likelihood estimator for Poisson regression with a misclassified binary exposure. In simulations, we observe that the MLE, on average, corrects for bias induced by the misclassified exposure in the naive analysis and reduces variance induced by deleting unqueried data in the complete case analysis. In our analysis of the Piedmont Triad, tracts with access to healthy foods had higher diabetes prevalence, on average, than those without. This difference was more pronounced for metropolitan tracts.
@@
We are potentially limited by our choice to assume a parametric logistic regression model for the exposure error mechanism. However, this choice has been shown to be fairly robust to misspecification (Lotspeich et al., 2022). Still, nonparametric or semiparametric error models are promising. Results from the case study are limited by the nature of the available data. Census tracts do not always fairly describe neighborhoods, but they were the most granular level of available diabetes data. Additionally, the model used to generate the small area diabetes estimates incorporates many socioeconomic factors that may have been relevant to our analysis but could not be included because of “double dipping.”
There are many interesting avenues of further exploration. It may be informative to explore spatial relationships among adjacent tracts, vary the parametric form of the outcome model, or modify the MLE to handle other types of data. The MLE’s efficiency could be further improved by a more strategic query design. Targeted strategies for selecting the most informative observations are common in the two-phase design literature.
@@
R package:
R package possum implementing the MLE (GNU zipped tar file)
@@
R scripts for the simulation studies and Piedmont Triad analysis (.zip file)

### v2

@@ 4.1 Data and Objectives
We investigated the associations between having access to healthy foods and diabetes prevalence across $N=387$ census tracts in the Piedmont Triad region of northwestern North Carolina. A census tract is a small, relatively permanent statistical subdivision of a county or statistically equivalent entity (United States Census Bureau, 2022). Approximately 1.7 million people live in the Piedmont Triad, and the region is known to have excellent transportation infrastructure. The model of interest is defined as $\log\left\{\textrm{E}_{\boldsymbol{\beta}}(Y\mid X,Z)\right\}=\beta_{0}+\beta_{1}X+\beta_{2}Z+\beta_{3}X\times Z+\log(O)$, where $Y$ is the number of diabetes cases, $O$ is the population (used as an offset), $X$ is an indicator of having access to healthy foods within a given radius, and $Z$ is an indicator of metropolitan status. We calculated the Haversine and route-based distances from the population-weighted centroid, according to the 2020 U.S. Census, to the nearest healthy food retailer. Using one-mile and half-mile radii, we discretized these distances to create our exposures $X$ and $X^{*}$. For example, $X=1$ if a census tract had at least one healthy food retailer within the threshold and =0=0 otherwise.
Queried data (i.e., with both $X$ and $X^{*}$) were available for all census tracts. However, to demonstrate the two-phase design, we assumed that only a subset had available route-based access data for the MLEs and complete case analyses. After running the gold standard analysis on the full data, we created a partially queried version wherein $X$ was only available for census tracts with $X^{*}=0$ (such that $X=0$ necessarily) and a subset of those with $X^{*}=1$. A random sample of $n=77$ tracts with $X^{*}=1$, targeting a 50/50 split between metropolitan/non-metropolitan, was chosen (Supplemental Figure S2). Only tracts with error-prone access (i.e., $X^{*}=1$) needed to be validated, since those with $X^{*}=0$ must have $X=0$ and do not inform the misclassification mechanism. The naive analysis used the error-prone access variable on the entire sample, and the complete case analysis used error-free access variable on the subset of queried data. The MLEs used the error-prone and error-free access indicators when available and just the error-prone version otherwise.
@@ 4.2 Diabetes Prevalence and Healthy Food Access
At both half- and one-mile radii, most tracts ($n=353$ $91\%$] and $n=265$ $68\%$], respectively) did not have access to healthy foods (Figure 1). Larger, rural tracts tended not to have access. Smaller, urban tracts tended to have access, or at least to be misclassified as such based on Haversine distances. Of the $192$ tracts with access to healthy foods based on one-mile Haversine distances, only $122$ actually had it based on route-based ones ($PPV=64\%$). As expected, the smaller threshold saw more misclassification ($PPV=40\%$).
@@
[…]
At either threshold, diabetes prevalence in neighborhoods with access to healthy foods was expected to be higher than in those without (Figure 2). This trend seems somewhat counterintuitive, but the literature does not agree on a clear pattern linking food access and diabetes (e.g., Gucciardi et al., 2014 versus Flint et al., 2020). Potential confounding (e.g., by socioeconomic status) or uncertainty around the outcome data (small area estimates) may play a role in this debate.
@@
The observed effect direction held at both distance thresholds and across all analysis methods. For example, based on the gold standard, diabetes prevalence in metropolitan tracts with access to healthy foods at one mile was expected to be 18% higher than in those without ($\widehat{\textrm{PR}}_{metro}=1.18$; 95% CI: $1.16$, $1.21$); among non-metropolitan tracts, those with access were expected to have 9% higher prevalence ($\widehat{\textrm{PR}}_{non}=1.09$; 95% CI: $1.06$, $1.11$).
The difference between tracts with and without access was more pronounced in metropolitan areas, according to the gold standard analyses. At both thresholds, the MLE also captured this difference, while the complete case analysis only did so at the one-half mile radius, and the naive analysis did not catch it at either threshold. The MLEs actually exaggerated this difference at the one-mile threshold but otherwise closely aligned with the gold standard analysis. Its 95% confidence intervals were slightly narrower than the complete case. For more details and further sensitivity analysis, see Web Appendix C.
@@ 5 Conclusion
We introduced a novel maximum likelihood estimator for Poisson regression with a misclassified binary exposure. In simulations, the MLE, on average, corrected for bias induced by the misclassified exposure in the naive analysis and reduced variance induced by deleting unqueried data in the complete case analysis. In our analysis of the Piedmont Triad, census tracts with access to healthy foods had higher diabetes prevalence, on average, than those without. This difference was more pronounced for metropolitan tracts.
@@
We are potentially limited by our choice to assume a parametric logistic regression model for the exposure error mechanism. However, this model has been shown to be fairly robust to misspecification (Lotspeich et al., 2022). Still, nonparametric or semiparametric error models are promising. Results from the case study are limited by the nature of the available data. Census tracts do not always fairly describe neighborhoods, but they were the most granular level of available diabetes data. Other studies (e.g., Rummo et al., 2021) consider food access at the individual level. If we had individual-level data, our proposed methods would still apply, simply omitting the model offset. Additionally, the model used to generate the small area diabetes estimates incorporates many socioeconomic factors that may have been relevant to our analysis but could not be included due to “double dipping.”
In the future, it may be informative to explore spatial relationships between adjacent tracts, vary the outcome model, or modify the MLE to handle other data types. The MLE’s efficiency could be further improved by a more strategic query design. Targeted strategies to select the most informative observations are common in the two-phase design literature.
(Flint et al., 2020)
@@
R package possum implementing the MLE (GNU zipped tar file) and R scripts for the simulation studies and Piedmont Triad analysis (.zip file)